OpenStudy (icalibear):

What is the simplified form of the expression -(10)^-1?

OpenStudy (cwrw238):

this = -1/10

OpenStudy (icalibear):

how did you get it?

OpenStudy (cwrw238):

the negative power = 1/ x another examples would be x^-2 = 1 / x^2 1/3^-1 = 1 / 3^1 = 1/3

OpenStudy (icalibear):

for that first example the simplified form would be 1/x^2 or 1/2?